A 20-cm-long, 190 g rod is pivoted at one end. A 19 g ball of clay is stuck on the other end.

What is the period if the rod and clay swing as a pendulum?
Answer #1

Length of the rod, L= 20cm=0.2m

Mass of the rod, M= 190g= 0.19kg

Mass of ball, m= 19g=0.019kg

Let Time period of the swing be "T".

Let \omega be the angular frequency of the swing, \theta be the angular frequency, then we can write the equation of motion of the rod about the pivot as,

(\frac{ML}{2}+mL)gsin \theta= -I\frac{d^2\theta}{dt^2}......................(1)

where g is the acceleration due to gravity, I be the total moment of inertia of the rod about the pivot and \theta be the small angular displacement of the rod.  

Then we can write moment of inertia of the rod as,

+ mL

Using above equation in equation 1,

  (\frac{ML}{2}+mL)gsin \theta=-(\frac{ML^2}{3}+mL^2) \frac{d^2\theta}{dt^2}

(\frac{M}{2}+m)Lgsin \theta=-(\frac{M}{3}+m) L^2\frac{d^2\theta}{dt^2}

Using small angle approximation: sin \theta\approx \theta , we get

(\frac{M+2m}{2})g\theta=-(\frac{M+3m}{3}) L\frac{d^2\theta}{dt^2}

\frac{d^2\theta}{dt^2}= -\frac{3(M+2m)g}{2L(M+3m)}\theta

Now comparing above equation with standard equation of motion, \frac{d^2\theta}{dt^2}= -\omega^2 \theta

We get,

\omega= \sqrt{\frac{3(M+2m)g}{2L(M+3m)}}

Using given values in above equation,

\omega= \sqrt{\frac{3[0.19+2(0.019)]9.8}{2\times 0.2[0.19+3(0.019)]}}

\omega= \sqrt{\frac{6.70}{0.0988}}= 8.23 rad/s

Then Time Period of the swing will be

T= \frac{ 2 \pi }{\omega}

Using the value of angular frequency in above equation,

T= \frac{ 2 \times 3.14 }{8.23}= 0.763s(ANS)

T= \frac{2.512 \times 0.247}{3\times 0.228 \times 9.8}=0.092s(ANS)